OpenStudy (saucvi):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of one half.

OpenStudy (saucvi):

A. y=1/2x+5 B. y=1/2x-3 C. y=1/2x-5 D. =1/2x+3

Vocaloid (vocaloid):

point-slope form: y - y1 = m(x-x1) plug in slope (1/2) for m, 2 for x1, and -4 for y1, then re-arrange and solve

OpenStudy (saucvi):

How do I plug in slope? I understand none of this ): @Vocaloid

Vocaloid (vocaloid):

replace m with 1/2 replace x1 with 2 replace y1 iwth -4

OpenStudy (saucvi):

@Vocaloid Is it y=1/2x-3

Vocaloid (vocaloid):

no, can you show me how you arrived at that answer?

OpenStudy (saucvi):

I plugged in the numbers and got y-(-4) = 1/2(x-2). Explaining it now I realize I don't even know how I got my answer. I think I rearranged wrong and miscalculated somewhere

OpenStudy (saucvi):

@vocaloid

Vocaloid (vocaloid):

yes, your plugging in is right but you may have made a mistake in the rearranging y-(-4) = 1/2(x-2) y+ 4 = 1/2(x) - 1 y = (1/2)x - 5

OpenStudy (saucvi):

Why did -4 become positive and where did the -1 come from? @vocaloid

Vocaloid (vocaloid):

4 becomes positive because y-(-4) = y + 4 (remember that two negative signs together become positive) 1/2(x-2) = 1/2(x) - 1 because you distribute the 1/2 1/2(x-2) = (1/2)(x) - (1/2)(2) = (1/2)x - 1

OpenStudy (saucvi):

Ohh okay. Thank you so much!